OpenStudy (anonymous):

how to determine: 1) mass of BaCl2.XH2O 2) mass of barium chloride anhydrous 3) mass of water in BaCl2.XH2O 4) Percent composition of water in BaCl2.xH2O 5) Formula of barium chloride hydrate

OpenStudy (anonymous):

THE INFORMATION mass of empty crucible = 56.5622g mass of crucible + hydrate before heating = 57.8310g mass of crucible + hydrate after heating (a) first reading = 57.6398g (b) second reading = 57.6374g (c) third reading = 57.6362g

OpenStudy (anonymous):

@gazed @JoannaBlackwelder

OpenStudy (joannablackwelder):

Any ideas on how to find the mass of the hydrate?

OpenStudy (anonymous):

i was thinking about "minus" the "before heating" and "the empty crucible."

OpenStudy (joannablackwelder):

Looks good to me :)

OpenStudy (anonymous):

then, how to get the mass of barium chloride anhydrous?

OpenStudy (joannablackwelder):

I would average the three after heating masses and then "minus" again from the "empty crucible"

OpenStudy (anonymous):

so i should find the "average first"?

OpenStudy (joannablackwelder):

Yeah, that's what I would do :)

OpenStudy (anonymous):

so, for "mass of water" = "average after heat" - "before heat" is it correct?

OpenStudy (joannablackwelder):

"before heat"-"average after heat"

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

i have the formula percentage composition of water percentage of water = mass of water loss / mass of sample x 100

OpenStudy (anonymous):

so is it... mass of BaCl2.XH2O / mass of water in BaCl2.XH2O x 100 ??

OpenStudy (joannablackwelder):

I would do percentage of water = mass of water/mass of hydrate x100

OpenStudy (anonymous):

Okay I will do like that, so how can i find the formula for the last question?

OpenStudy (joannablackwelder):

I think we need to use the molar masses of BaCl2 and H2O to find the moles of each. Then use a ratio to find the composition.

OpenStudy (anonymous):

i still cant understand the last part

OpenStudy (joannablackwelder):

Can you find the moles of BaCl2?

OpenStudy (joannablackwelder):

*Hint: use molar mass and mass of anhydrous

OpenStudy (anonymous):

1) 57.8310g - 56.5622g = 1.2688 g hydrate 2) 57.6362g - 56.5622g = 1.07400 g anhydride 3) 57.8310g - 57.6362g = 0.1948 g water 4) (0.1948 g water) / (1.2688 g hydrate) = 0.1535 = 15.35% water 5) 15.35% water = 84.65% BaCl2 (208.2330 g/mol BaCl2) / (0.8465) = 245.99 g/mol hydrate ((245.99 g/mol) - (208.2330 g/mol BaCl2)) / (18.01532 g H2O/mol) = 2.096 Round to the nearest whole number (2) to find the molar ratio of water to BaCl2: BaCl2·2H2O